OpenStudy (anonymous):

Lyle has a ten-cup box of cereal. If one serving is 3/4 cup, how many servings are in the box?

OpenStudy (amistre64):

10/(3/4) = 40/3

OpenStudy (anonymous):

13.333

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{T}}\color{orange}{\normalsize\text{h}}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{n}}\color{blue}{\normalsize\text{k}}\color{purple}{\normalsize\text{ }}\color{purple}{\normalsize\text{U}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{P}}\color{#9c9a2e}{\normalsize\text{e}}\color{green}{\normalsize\text{o}}\color{blue}{\normalsize\text{p}}\color{purple}{\normalsize\text{l}}\color{purple}{\normalsize\text{e}}\color{red}{\normalsize\text{!}}\color{orange}{\normalsize\text{!}}\color{#9c9a2e}{\normalsize\text{}}\end{array}

OpenStudy (amistre64):

nice color :)

OpenStudy (anonymous):

thank u